# Are foundation models for computer vision good conformal predictors?

### v2

@@ 4.2.4 Effect on few-shot adapted models
Adapting zero-shot CLIP models for downstream tasks in a few-shot labeled regime is becoming increasingly popular in VLMs. These strategies can be mainly categorized into Prompt Learning [70, 17, 69], which optimize the set of text prompts given to the text encoder, and Adapters [67, 30, 23, 53, 70, 17], where only a limited set of learnable parameters atop embedding representations is updated. Thus, questioning whether adapting these models hinges the performance of CP methods is of paramount importance, as it addresses a foundational aspect of effectively integrating uncertainty quantification with modern architectures. To assess the impact of CLIP adaptation in conformal prediction, we resort to representative methods of the few-shot adaptation families presented above, and follow standard adaptation and evaluation protocols in the literature [53]: (i) for out-of distribution (OOD), we adapt CLIP on few-shot samples ($M=16$) from ImageNet, and evaluate on ImageNet and its variants, and (ii) for in distrbution (ID) we adapt on few-shot samples and evaluate on the validation set from the same dataset.
@@
CLAP [53] 2.13 3.52 2.57 0.088 0.060 0.065
CoOp [70] 2.07 2.87 2.47 0.083 0.059 0.067
KgCoOp [17] 2.10 3.11 2.53 0.085 0.059 0.065
@@
CLAP [53] 7.55 17.41 9.89 0.093 0.083 0.091
CoOp [70] 7.86 16.35 10.03 0.093 0.085 0.095
KgCoOp [17] 7.54 16.68 9.79 0.094 0.085 0.095
Table 3: Few-shot adaptation (16-shots). Set size and CovGap for CLIP (ViT-B backbone) on in-distribution (average over 11 datasets) and out-of-distribution (average over ImageNet versions). Results with 4-shots are provided in Appendix, Table 10.
Few-shot VLM adaptation render pre-trained models better conformalized in the ID scenario (Table 3). First, we observe that, across both ID and OOD scenarios, both Adapters and Prompt Learning yield smaller conformal set sizes and higher Coverage gap than zero-shot CLIP, regardless of the CP method, with Prompt Learning yielding slightly better performances. In contrast, on the OOD scenario, only APS consistently enhances the set efficiency of the ZS model, with all the few-shot adaptation methods obtaining scarce coverage gap improvements across CP methods. These observations are related to the findings from the previous section, which suggested that better-calibrated models lead to larger conformal sets. Indeed, recent evidence [36] demonstrated that few-shot CLIP adaptation methods deteriorate the confidence estimates compared to ZS predictions, which, following our observations, should result in smaller conformal sets, as validated in Table 3. Further details are provided in Appendix G.
@@ 5 Conclusion
Acknowledgments
@@
Supplementary Material
A Metrics
Given a pre-specified coverge probability $1-\alpha$, a test set $\mathcal{D}_{\text{test}}$ and a method for predicting conformal sets $\hat{C}(\cdot)$, we resort to the following metrics.
Set $\text{Set size}(\mathcal{D}_{\text{test}})=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{\mathbf{x}\in\mathcal{D}_{\text{test}}}|\hat{C}(\mathbf{x})|$ (7)
Where Equation 7 is the average predicted set size across the test set (sometimes refered to as efficiency), Equation 8 is the empirical marginal coverage, Equation 9 represents the average gap between the empirical coverage and the class-conditioned coverage, and Equation 10 is the min class-conditioned coverage (MCCC).
B Models Used and Implementation Details
Foundation models.
@@ B Models Used and Implementation Details
We employ three different, and diverse foundation models for vision tasks, which differ in the number of modalities for training (i.e., vision vs vision-language), strategy used for training (i.e., contrastive vs self-supervised learning), and backbones used (i.e., vision transformers vs convolutional neural networks). More conctretely, the DINO [8, 40] models employed in this work are ViT-based models trained in a self-supervised manner. CLIP [46] is a vision-language model, which trains a vision and a text encoder in a contrastive fashion. Last, VICReg [3] are CNNs trained in a self-supervised manner using contrastive learning.
Implementation details.
@@
All linear probing heads are trained with a cross-entropy loss with the Adam [25] optimizer and a learning rate of $10^{-4}$, whose patience is set to $10$. The implementation of the training is based on PyTorch [42], and the conformal predictions methods on TorchCP [63]. For RAPS, we used $k_{\text{RAPS}}=2$ and $\lambda_{\text{RAPS}}=0.1$ as hyperparameters. Across all experiments, we use $\alpha=0.1$, except for CIFAR-10 for which we use $\alpha=0.05$.
C Relationship Between Model Performance and CP Metrics
@@
Table 4, Table 5, and Table 6 report the numerical values corresponding to Fig. 1, showcasing the relationship between the linear probing accuracy of each model, and the CP performance in terms of set size and coverage gap for ImageNet, CIFAR-10, and CIFAR-100, respectively.
D Model performance for the optimal temperature
Fig. 8, 9, 10, 11, 12, and 13 showcase the impact of temperature $T$ on the ECE, set size, minimum class-conditional coverage, and coverage gap across all foundation models. The observations on these figures strongly align with the findings on the main paper as, regardless of the model, APS yields optimal coverage gap close to the optimal temperature point. Indeed, looking at the different models (i.e., DINO-based, VICReg and CLIP), we can observe that the behaviour of RAPS (in terms of coverage gap), strongly varies across models. Additionally, Fig. 14 shows the evolution of $q_{\alpha}$, the threshold on the conformal scores $s$ before and after calibration. When applying temperature scaling, the distribution of softmax across classes approaches a uniform distribution, lowering the scores for the most likely classes. Intuitively, this means that more classes will need to be included in the set to ensure coverage, leading to a decrease in threhsold $q_{\alpha}$. This decrease is consistent with the observed increase in set size, which can be seen in Fig. 15, showing the difference in set size when applying temperature scaling for APS on CLIP. Note that this trend holds across models, as a similar behaviour for DINO is observed in Fig. 7 in the main paper. Note that both Fig. 7 and 15 only show cases where the set sizes are different between the calibrated and uncalibrated model.
@@ D Model performance for the optimal temperature
Furthermore, we report in Table 7 a more complete version of the results presented in Table 2, containing all explored models.
E Conformal set size analysis
@@
We show the difference between the conformal set size for APS when applied to $\text{ViT}_{\text{ImageNet}}$ and $\text{ViT}_{\text{DINO-S}}$ in Fig. 20 and for $\text{ViT}_{\text{ImageNet}}$ and $\text{ViT}_{\text{MetaCLIP}}$ in Fig. 19, similarly to what is shown in Fig. 4.
F Domain Shift
@@
Table 8 reports the numerical values for Fig. 4, showcasing APS’s strong performance in terms of marginal and conditional coverage under distribution shift, at the cost of decreasing the set efficiency. Fig. 16, 17, and 18 depict the distribution of class-conditional coverage for DINOv2-B, VICReg (ResNet 50x2), and CLIP, respectively. The kernel density estimate plots show the class-conditional coverage for a distribution shift with ImageNet-A (top), ImageNet-R (middle), and ImageNet-V2 (bottom), for APS (left) and RAPS (right). These results clearly demonstrate the stronger resistance to distribution shift for APS compared to RAPS, even when both methods show a minimum class-conditional coverage of 0.
G Few-Shot Adaptation

### v3

@@ 4.2.4 Effect on few-shot adapted models
Adapting zero-shot CLIP models for downstream tasks in a few-shot labeled regime is becoming increasingly popular in VLMs. These strategies can be mainly categorized into Prompt Learning (Zhou et al., 2022b; Hantao Yao, 2023; Zhou et al., 2022a), which optimize the set of text prompts given to the text encoder, and Adapters (Yu et al., 2023; Lin et al., 2023; Huang et al., 2024; Silva-Rodríguez et al., 2024; Zhou et al., 2022b; Hantao Yao, 2023), where only a limited set of learnable parameters atop embedding representations is updated. Thus, questioning whether adapting these models hinges the performance of CP methods is of paramount importance, as it addresses a foundational aspect of effectively integrating uncertainty quantification with modern architectures. To assess the impact of CLIP adaptation in conformal prediction, we resort to representative methods of the few-shot adaptation families presented above, and follow standard adaptation and evaluation protocols in the literature (Silva-Rodríguez et al., 2024): (i) for out-of distribution (OOD), we adapt CLIP on few-shot samples ($M=16$) from ImageNet, and evaluate on ImageNet and its variants, and (ii) for in distribution (ID) we adapt on few-shot samples and evaluate on the validation set from the same dataset.
Table 3: Few-shot adaptation (16-shots). Set size and CovGap for CLIP (ViT-B backbone) on in-distribution (average over 11 datasets) and out-of-distribution (average over ImageNet versions). Additional results are presented in Appendix Section F.
@@
CLAP (Silva-Rodríguez et al., 2024) 2.13 3.52 2.57 0.088 0.060 0.065
CoOp (Zhou et al., 2022b) 2.07 2.87 2.47 0.083 0.059 0.067
KgCoOp (Hantao Yao, 2023) 2.10 3.11 2.53 0.085 0.059 0.065
@@
CLAP (Silva-Rodríguez et al., 2024) 7.55 17.41 9.89 0.093 0.083 0.091
CoOp (Zhou et al., 2022b) 7.86 16.35 10.03 0.093 0.085 0.095
KgCoOp (Hantao Yao, 2023) 7.54 16.68 9.79 0.094 0.085 0.095
Few-shot VLM adaptation of pre-trained models lead to lower set sizes and coverage gaps in the ID scenario (Table 3). First, we observe that, across both ID and OOD scenarios, both Adapters (ZSLP and CLAP) and Prompt Learning (CoOp and KgCoOp) yield smaller conformal set sizes and higher coverage gap than zero-shot CLIP, regardless of the CP method, with Prompt Learning yielding slightly better performances. In contrast, on the OOD scenario, only APS consistently enhances the set efficiency of the ZS model, with all the few-shot adaptation methods obtaining scarce coverage gap improvements across CP methods. These observations are related to the findings from the previous section, which suggested that better-calibrated models lead to larger conformal sets. Indeed, recent evidence (Murugesan et al., 2024a) demonstrated that few-shot CLIP adaptation methods deteriorate the confidence estimates compared to ZS predictions, which, following our observations, should result in smaller conformal sets, as validated in Table 3. Further details are provided in Appendix Section F.
@@ 5 Conclusion
The final decision ultimately hinges on the requirements of each task. In certain fields, e.g., medical diagnosis, it may be preferable to maximize the conditional coverage across classes, even if it increases the conformal set size, as this minimizes the risk of critical errors. In contrast, in domains where errors are less consequential, maintaining smaller set sizes might be preferable to streamline decision-making and reduce computational demands. Thus, the choice depends on balancing the trade-offs between accuracy, usability, and the specific demands of the application.
Acknowledgement
@@
Appendix A Models Used and Implementation Details
@@
Foundation models. We use 17 foundation models for vision tasks from three categories. They differ in the number of modalities for training (i.e., vision vs vision-language), strategy used for training (i.e., contrastive vs self-supervised learning), and backbones used (i.e., vision transformers vs convolutional neural networks). More conctretely, the DINO (Caron et al., 2021; Oquab et al., 2024) models employed in this work are ViT-based models trained in a self-supervised manner. CLIP (Radford et al., 2021) is a vision-language model, which trains a vision and a text encoder in a contrastive fashion. Last, VICReg (Bardes et al., 2022) are CNNs trained in a self-supervised manner using contrastive learning. Table 4 summarizes the models used for this study.
@@
Implementation details. All linear probing heads are trained with a cross-entropy loss with the Adam (Kingma & Ba, 2017) optimizer and a learning rate of $10^{-4}$, whose patience is set to $10$. The implementation of the training is based on PyTorch (Paszke et al., 2019), and the conformal predictions methods on TorchCP (Wei & Huang, 2024). For RAPS, we used $k_{\text{RAPS}}=2$ and $\lambda_{\text{RAPS}}=0.1$ as hyperparameters. Across all experiments, we use $\alpha=0.1$, except for CIFAR-10 for which we use $\alpha=0.05$.
Appendix B Relationship Between Model Performance and CP Metrics
@@
Table 5, Table 6, and Table 7 report the numerical values corresponding to Figure 1, showcasing the relationship between the linear probing accuracy of each model, and the CP performance in terms of set size and coverage gap for CIFAR-10, and CIFAR-100, and ImageNet respectively.
Appendix C Model Calibration
C.1 Temperature Scaling
@@
We report in Table 8 a more complete version of the results presented in Table 2, containing all explored models.
Additionally, Figures 10, 11, 12, 13, 14, 15, 16, 17, 18 showcase the impact of temperature $T$ on the ECE, set size, minimum class-conditional coverage, and coverage gap across all foundation models. The observations on these figures strongly align with the findings on the main paper as, regardless of the model, APS yields optimal coverage gap close to the optimal temperature point. Indeed, looking at the different models (i.e., DINO-based, VICReg and CLIP), we can observe that the behaviour of RAPS (in terms of coverage gap), strongly varies across models. Additionally, Figure 8 shows the evolution of $q_{\alpha}$, the threshold on the conformal scores $s$ before and after calibration. When applying temperature scaling, the distribution of softmax across classes approaches a uniform distribution, lowering the scores for the most likely classes. Intuitively, this means that more classes will need to be included in the set to ensure coverage, leading to a decrease in threhsold $q_{\alpha}$. This decrease is consistent with the observed increase in set size, which can be seen in Fig. 9, showing the difference in set size when applying temperature scaling for APS on CLIP. Note that this trend holds across models, as a similar behaviour for DINO is observed in Fig. 6 in the main paper. Note that both Fig. 6 and 9 only show cases where the set sizes are different between the calibrated and uncalibrated model.
C.2 Histogram Binning
In Table 9, we explore histogram binning as an alternative method (other than temperature scaling) for model calibration. This method sorts predictions into a predefined number of bins ($100$ in our case) and adjusts the predictions for the confidence score to match the correct frequency. This is learned on a separate set and applied on a test set. In particular, with this approach as calibration technique we observed an increase in set size, similar to TS, with a slight decrease in MCCC, aligning with our observations on the effects of calibration on conformalization.
Appendix D Conformal set size analysis
@@
We show the difference between the conformal set size for APS when applied to $\text{ViT}_{\text{ImageNet}}$ and $\text{ViT}_{\text{DINO-S}}$ and for $\text{ViT}_{\text{ImageNet}}$ and $\text{ViT}_{\text{MetaCLIP}}$ in Figure 19, similarly to what is shown in Figure 5.
Appendix E Domain Shift
@@
Table 10 reports the numerical values for Fig. 5, showcasing APS’s strong performance in terms of marginal and conditional coverage under distribution shift, at the cost of decreasing the set efficiency. Figure 21 depict the distribution of class-conditional coverage for DINOv2-B, VICReg (RN 50x2), and CLIP (ViT-B). The kernel density estimate plots show the class-conditional coverage for a distribution shift with ImageNet-A (top), ImageNet-R (middle), and ImageNet-V2 (bottom), for APS (left) and RAPS (right). These results clearly demonstrate the stronger resistance to distribution shift for APS compared to RAPS, even when both methods show a minimum class-conditional coverage of 0.
Appendix F Few-Shot Adaptation